# Targeted Deep Learning System Boundary Testing

### v2

@@ 4.7. Qualitative Analysis
While metrics and quantitative analysis allows for comparability of results, they often do not communicate the full picture adequately, especially with respect to image realism (Lambertenghi and Stocco, 2024). To showcase the performance differences between Mimicry and the baseline methods, we perform a qualitative analysis of the produced outputs by manually analyzing a few meaningful examples.
The first interesting aspect relates to the Laplacian Variance of image differences seen in (missing) 10. This numerical measure it is not widely used and therefore needs more explanation and positioning. Especially in comparing Mimicry to Sinvad it is useful, while it fails for the comparison with DeepJanus. As described earlier with this measure the type of change in the image across the optimization process is quantified. Low Laplacian variance in the image differences here means that the image gets gradually blurred, not producing functional differences in the output candidate (10(b) & 10(e)). This can easily be seen in the examples for Sinvad, where with more complex data, this phenomenon gets more prominent. In contrast, Mimicry does not blur the original seed images, rather produces functionally different outputs, even if those outputs may no longer convey clear class information to the human observer.
@@ 4.8.1. Internal validity
All experiments of Mimicry and the baselines have been conducted on the same computational budget, based on SUT predictions. Additionally, regarding the StyleGAN models, we utilized pre-trained models available from the literature (Sauer et al., 2022; Karras et al., 2020b). When not available, we trained the StyleGAN models using the scripts available in the replication package of the original paper (Karras et al., 2020b), as it is difficult to envision less threat-prone approaches. In the human evaluation study, we included attention questions to ensure that participants were engaging truthfully (Sorokin and Forsyth, 2008). Since both DeepJanus and Sinvad require input images, rather than generating inputs like Mimicry, the input selection process was randomized. This approach aims to reduce bias and ensures a more fair comparison with Mimicry.
@@ 4.8.2. External validity
The limited number of ML systems included in our evaluation poses a threat to the generalizability of our results. We addressed this issue by incorporating a variety of datasets with increasing complexity and high-performing models from related literature. We demonstrated the usefulness of the StyleGAN architecture (Karras et al., 2020b; Sauer et al., 2022); however, other style-based architectures (Karras et al., 2019b, 2021) may also yield promising results.
@@ 5. Discussion
Generating well-defined, unambiguous boundary inputs for high-resolution datasets remains an open challenge. Mimicry produces inputs with high validity rates, provided the generated images depict a class that is perceptible to human observers. In terms of boundary preservation, Mimicry generally performs well except in the case of ImageNet. Unlike Sinvad, which mainly blurs existing dataset images, Mimicry generates functionally novel inputs that may sometimes appear ambiguous to humans. For ImageNet, the relatively high resolution ($128\times 128$) allows the blur introduced by Sinvad to retain enough visual cues for human evaluators to identify the original class. In contrast, Mimicry may generate objects that are less recognizable or entirely unfamiliar, given the feature mixing between two classes. This highlights an open challenge: as dataset complexity increases, assessing the validity of generated test cases becomes more difficult for human evaluators, especially for classes with low to no semantic affinity. In such contexts, alternative evaluation methods such as using large language models as judges (Gu et al., 2025) might be of interest to corroborate the human assessment.
@@ 6.1. Model-based Input Manipulation
Model Input Manipulation (MIM) techniques leverage a model of the input domain to generate test inputs, similar to conventional model-driven engineering practices that uphold compliance with domain-specific constraints (Abdessalem et al., 2018; Ben Abdessalem et al., 2016, 2018; Gambi et al., 2019; Riccio and Tonella, 2020; Neelofar and Aleti, 2024).
The manipulation occurs on the model, which is subsequently reconverted to the original format (Larman et al., 1998). MIM techniques operate within a restricted input space, specifically the control parameters of the model representation. These techniques enhance the realism of the produced outputs by implementing appropriate model constraints.
Several search-based MIM approaches have been applied to DL-based image classifiers. DeepHyperion (Zohdinasab et al., 2021) uses the MAP-Elites Illumination Search algorithm (Mouret and Clune, 2015) to explore the feature space of the input domain and identify misbehavior-inducing features. DeepMetis (Riccio et al., 2021) a MIM approach that generates inputs that behave correctly on original DL models and misbehave on mutants obtained through injection of realistic faults (Humbatova et al., 2020), which can be useful to enhance the mutation killing ability of a test set. DeepJanus (Riccio and Tonella, 2020) is the MIM approach most related to this work since it performs boundary testing of DL systems. Therefore, we performed an explicit empirical comparison with the DeepJanus approach in this work.
However, a significant limitation of MIM approaches is their reliance on the availability of a high-quality model representation for the specific input domain, which is manually crafted (Riccio and Tonella, 2023). Unlike MIM techniques, Mimicry leverages a generative network to learn the distribution of the input domain. This approach is largely automated and requires no labeling or other cost, except for hyperparameter tuning. This characteristics of Mimicry broadens its applicability across various domains.
@@ 6.2. Raw Input Manipulation
Raw input manipulation (RIM) techniques involve modifying an image’s original pixel space to create a new input by perturbing the pixel values. RIM techniques aim to produce minimal, often imperceptible changes to original to trigger misbehavior in the DL system (Liu et al., 2022; Zhang et al., 2020a; Zohdinasab et al., 2021; Kurakin et al., 2018; Croce and Hein, 2020). These methods do not focus on boundary analysis and target different aspects of testing, such as data augmentation or adversarial attacks, which are not directly aligned with our goal. Our method is a functional test generator, differing from adversarial testing in both goals and techniques. Functional testing creates new, valid, in-distribution inputs to evaluate a DNN’s generalization. In contrast, adversarial testing adds minor perturbations to original inputs to test robustness (Croce and Hein, 2020). Given these distinct objectives and methods, direct comparisons are inappropriate. However, for completeness, we describe the main propositions next.
DeepXplore (Pei et al., 2017) employs various techniques, including occlusion, light manipulation, and blackout to cause misbehavior. These perturbations are intended to improve neuron coverage within the DL system. DLFuzz (Guo et al., 2018) introduces noise to the seed image to increase the likelihood of system misbehavior. DLFuzz generates adversarial inputs for DL systems without relying on cross-referencing other similar DL systems or manual labeling. DeepTest (Tian et al., 2018) alters the images using synthetic affine transformation from the computer vision domain, such as blurring and brightness adjustments, to create simulated rain/fog effects.
RIM techniques are limited to modifying existing inputs and they cannot thoroughly explore the input domain and its boundaries, while generative DL models can sample novel inputs from the data distribution. Moreover, the manipulated images might not always represent real-world functional inputs, e.g., images with artificial artifacts at the corners or unnatural lighting conditions generated by DeepXplore. Consequently, such techniques are more suitable for security and robustness testing rather than for functional testing (Riccio and Tonella, 2023).
@@ 6.3. Latent Space Manipulation
Latent space manipulation techniques generate new inputs by learning and reconstructing the underlying distribution of the input data. The most commonly used techniques are Variational Autoencoders (VAE) (Kingma, 2013) and Generative Adversarial Networks (GAN) (Goodfellow et al., 2020).
Sinvad (Kang et al., 2020, 2024) constructs the input space using VAE and navigates the latent space by adding a random value sampled from a normal distribution to a single element of the latent vector. Sinvad aims to explore the latent space by maximizing either the probability of misbehaviors, estimated from the softmax layer output, or by surprise coverage (Kim et al., 2019).
The Feature Perturbations technique (Dunn et al., 2021, 2020) involves injecting perturbations into the output of the generative model’s first layers, which represent high-level features of images. These perturbations can affect various characteristics of the image, such as shape, location, texture, or color. DeepRoad (Zhang et al., 2018) generates driving images using Generative Adversarial Networks (GANs) for image-to-image translation.
CIT4DNN (Dola et al., 2024) combines VAE and combinatorial testing (Cohen et al., 1997). This allows the systematic exploration and generation of diverse and infrequent input datasets. CIT4DNN partitions latent spaces to create test sets that contain a wide range of feature combinations and rare occurrences. A recently proposed technique, Instance Space Analysis, aims to pinpoint the critical features of test scenarios that impact the detection of unsafe behavior (Neelofar and Aleti, 2024).
@@ 7. Conclusion & Future Work
In this work, we present Mimicry, a tool for targeted boundary testing of ML classifiers by identifying inputs near decision boundaries. Our empirical analysis demonstrates that Mimicry outperforms existing methods such as DeepJanus (Riccio and Tonella, 2020) and Sinvad (Kang et al., 2020), particularly in complex data domains, by leveraging latent space manipulations and incorporating SUT behavior into the search process. Unlike DeepJanus, which relies on model representations of inputs, and Sinvad, which suffers from limited control over generative manipulations, Mimicry effectively balances control, fidelity, and performance in generating meaningful boundary test cases.
@@ 8. Data Availability
The experiment codebase, analysis scripts and all artifacts generated for this work can be found in the replication package (rep, 2025). Artifacts that were generated by other works are linked accordingly.

### v3

@@ 4.7. Qualitative Analysis
While metrics and quantitative analysis allows for comparability of results, they often do not communicate the full picture adequately, especially with respect to image realism (lambertenghi2024assessing). To showcase the performance differences between Mimicry and the baseline methods, we perform a qualitative analysis of the produced outputs by manually analyzing a few meaningful examples.
The first interesting aspect relates to the Laplacian Variance of image differences seen in () 10. This numerical measure it is not widely used and therefore needs more explanation and positioning. Especially in comparing Mimicry to Sinvad it is useful, while it fails for the comparison with DeepJanus. As described earlier with this measure the type of change in the image across the optimization process is quantified. Low Laplacian variance in the image differences here means that the image gets gradually blurred, not producing functional differences in the output candidate (10(b) & 10(e)). This can easily be seen in the examples for Sinvad, where with more complex data, this phenomenon gets more prominent. In contrast, Mimicry does not blur the original seed images, rather produces functionally different outputs, even if those outputs may no longer convey clear class information to the human observer.
@@ 4.8.1. Internal validity
All experiments of Mimicry and the baselines have been conducted on the same computational budget, based on SUT predictions. Additionally, regarding the StyleGAN models, we utilized pre-trained models available from the literature (sauer2022styleganXL; karras2020analyzing). When not available, we trained the StyleGAN models using the scripts available in the replication package of the original paper (karras2020analyzing), as it is difficult to envision less threat-prone approaches. In the human evaluation study, we included attention questions to ensure that participants were engaging truthfully (sorokin2008utility). Since both DeepJanus and Sinvad require input images, rather than generating inputs like Mimicry, the input selection process was randomized. This approach aims to reduce bias and ensures a more fair comparison with Mimicry.
@@ 4.8.2. External validity
The limited number of ML systems included in our evaluation poses a threat to the generalizability of our results. We addressed this issue by incorporating a variety of datasets with increasing complexity and high-performing models from related literature. We demonstrated the usefulness of the StyleGAN architecture (karras2020analyzing; sauer2022styleganXL); however, other style-based architectures (karras2019style; karras2021alias) may also yield promising results.
@@ 5. Discussion
Generating well-defined, unambiguous boundary inputs for high-resolution datasets remains an open challenge. Mimicry produces inputs with high validity rates, provided the generated images depict a class that is perceptible to human observers. In terms of boundary preservation, Mimicry generally performs well except in the case of ImageNet. Unlike Sinvad, which mainly blurs existing dataset images, Mimicry generates functionally novel inputs that may sometimes appear ambiguous to humans. For ImageNet, the relatively high resolution ($128\times 128$) allows the blur introduced by Sinvad to retain enough visual cues for human evaluators to identify the original class. In contrast, Mimicry may generate objects that are less recognizable or entirely unfamiliar, given the feature mixing between two classes. This highlights an open challenge: as dataset complexity increases, assessing the validity of generated test cases becomes more difficult for human evaluators, especially for classes with low to no semantic affinity. In such contexts, alternative evaluation methods such as using large language models as judges (gu2025surveyllmasajudge) might be of interest to corroborate the human assessment.
@@ 6.1. Model-based Input Manipulation
Model Input Manipulation (MIM) techniques leverage a model of the input domain to generate test inputs, similar to conventional model-driven engineering practices that uphold compliance with domain-specific constraints (Abdessalem-ASE18-1; Abdessalem-ASE18-2; Abdessalem-ICSE18; Gambi:2019:ATS:3293882.3330566; riccio2020deepjanus; isa).
The manipulation occurs on the model, which is subsequently reconverted to the original format (larman1998applying). MIM techniques operate within a restricted input space, specifically the control parameters of the model representation. These techniques enhance the realism of the produced outputs by implementing appropriate model constraints.
Several search-based MIM approaches have been applied to DL-based image classifiers. DeepHyperion (zohdinasab2021deephyperion) uses the MAP-Elites Illumination Search algorithm (DBLP:journals/corr/MouretC15) to explore the feature space of the input domain and identify misbehavior-inducing features. DeepMetis (2021-Riccio-ASE) a MIM approach that generates inputs that behave correctly on original DL models and misbehave on mutants obtained through injection of realistic faults (2020-Humbatova-ICSE), which can be useful to enhance the mutation killing ability of a test set. DeepJanus (riccio2020deepjanus) is the MIM approach most related to this work since it performs boundary testing of DL systems. Therefore, we performed an explicit empirical comparison with the DeepJanus approach in this work.
However, a significant limitation of MIM approaches is their reliance on the availability of a high-quality model representation for the specific input domain, which is manually crafted (2023-Riccio-ICSE). Unlike MIM techniques, Mimicry leverages a generative network to learn the distribution of the input domain. This approach is largely automated and requires no labeling or other cost, except for hyperparameter tuning. This characteristics of Mimicry broadens its applicability across various domains.
@@ 6.2. Raw Input Manipulation
Raw input manipulation (RIM) techniques involve modifying an image’s original pixel space to create a new input by perturbing the pixel values. RIM techniques aim to produce minimal, often imperceptible changes to original to trigger misbehavior in the DL system (liu2022deepboundary; zhang2020deepsearch; zohdinasab2021deephyperion; kurakin2018adversarial; croce2020minimally). These methods do not focus on boundary analysis and target different aspects of testing, such as data augmentation or adversarial attacks, which are not directly aligned with our goal. Our method is a functional test generator, differing from adversarial testing in both goals and techniques. Functional testing creates new, valid, in-distribution inputs to evaluate a DNN’s generalization. In contrast, adversarial testing adds minor perturbations to original inputs to test robustness (croce2020minimally). Given these distinct objectives and methods, direct comparisons are inappropriate. However, for completeness, we describe the main propositions next.
DeepXplore (pei2017deepxplore) employs various techniques, including occlusion, light manipulation, and blackout to cause misbehavior. These perturbations are intended to improve neuron coverage within the DL system. DLFuzz (guo2018dlfuzz) introduces noise to the seed image to increase the likelihood of system misbehavior. DLFuzz generates adversarial inputs for DL systems without relying on cross-referencing other similar DL systems or manual labeling. DeepTest (deeptest) alters the images using synthetic affine transformation from the computer vision domain, such as blurring and brightness adjustments, to create simulated rain/fog effects.
RIM techniques are limited to modifying existing inputs and they cannot thoroughly explore the input domain and its boundaries, while generative DL models can sample novel inputs from the data distribution. Moreover, the manipulated images might not always represent real-world functional inputs, e.g., images with artificial artifacts at the corners or unnatural lighting conditions generated by DeepXplore. Consequently, such techniques are more suitable for security and robustness testing rather than for functional testing (2023-Riccio-ICSE).
@@ 6.3. Latent Space Manipulation
Latent space manipulation techniques generate new inputs by learning and reconstructing the underlying distribution of the input data. The most commonly used techniques are Variational Autoencoders (VAE) (kingma2013auto) and Generative Adversarial Networks (GAN) (goodfellow2020generative).
Sinvad (kang2020sinvad; sinvad-tosem) constructs the input space using VAE and navigates the latent space by adding a random value sampled from a normal distribution to a single element of the latent vector. Sinvad aims to explore the latent space by maximizing either the probability of misbehaviors, estimated from the softmax layer output, or by surprise coverage (kim2019guiding).
The Feature Perturbations technique (dunn2021exposing; DBLP:journals/corr/abs-2001-11055) involves injecting perturbations into the output of the generative model’s first layers, which represent high-level features of images. These perturbations can affect various characteristics of the image, such as shape, location, texture, or color. DeepRoad (deeproad) generates driving images using Generative Adversarial Networks (GANs) for image-to-image translation.
CIT4DNN (dola2024cit4dnn) combines VAE and combinatorial testing (cit605761). This allows the systematic exploration and generation of diverse and infrequent input datasets. CIT4DNN partitions latent spaces to create test sets that contain a wide range of feature combinations and rare occurrences. A recently proposed technique, Instance Space Analysis, aims to pinpoint the critical features of test scenarios that impact the detection of unsafe behavior (isa).
@@ 7. Conclusion & Future Work
In this work, we present Mimicry, a tool for targeted boundary testing of ML classifiers by identifying inputs near decision boundaries. Our empirical analysis demonstrates that Mimicry outperforms existing methods such as DeepJanus (riccio2020deepjanus) and Sinvad (kang2020sinvad), particularly in complex data domains, by leveraging latent space manipulations and incorporating SUT behavior into the search process. Unlike DeepJanus, which relies on model representations of inputs, and Sinvad, which suffers from limited control over generative manipulations, Mimicry effectively balances control, fidelity, and performance in generating meaningful boundary test cases.
@@ 8. Data Availability
The experiment codebase, analysis scripts and all artifacts generated for this work can be found in the replication package (replication-package). Artifacts that were generated by other works are linked accordingly.